# Robust Bayesian Inference for Berkson and Classical Measurement Error Models

### v2

@@ 4.1 Target distribution & assumptions
The kernel $k$ is $\mathfrak{S}_{\mathcal{Z}}$-measurable and satisfies $|k|\leq 1$. Moreover, $\phi:\mathcal{Z}\rightarrow\mathcal{H}_{k}$ is such that $k((x,y),(x^{\prime},y^{\prime}))=\left<\phi(x,y),\phi(x^{\prime},y^{\prime})\right>_{\mathcal{H}_{k}}$.
@@
There is a continuous kernel $k_{X}$ on $\mathcal{X}$ and a kernel $k_{Y}$ on $\mathcal{Y}$ such that $|k_{X}|\leq 1$, $|k_{Y}|\leq 1$ and $k=k_{X}\otimes k_{Y}$.
As noted in Alquier and Gerber, (2024), from Szabó and Sriperumbudur, (2017) it follows that under Assumption 4.0.2, $k$ is characteristic if $k_{X}$ and $k_{Y}$ are characteristic, continuous, translation invariant and bounded. When $\mathcal{X}=\operatorname{\mathbb{R}}^{d_{x}}$ and $\mathcal{Y}=\operatorname{\mathbb{R}}^{d_{y}}$ examples of such kernels $k_{X}$ and $k_{Y}$ include the widely-used RBF, exponential kernel and Matérn kernels. Here we restrict our results to the case where $k_{X}$ is a translation invariant kernel:
@@
There exists a bounded, continuous, real-valued, positive definite function $\psi$ on $M=\mathbb{R}^{d_{x}}$ such that $k_{X}(x,x^{\prime})=\psi(x-x^{\prime})$.
While this assumption is not strictly necessary for deriving generalization error bounds similar to the ones that follow, it demonstrates the error decomposition relative to the MMD between the true ME distribution and the prior centering measure. This choice also makes intuitive sense for the case of additive ME, as in this work, since the distance between any two points $x+\nu_{1}$ and $x+\nu_{2}$ is characterized by the distance between $\nu_{1}$ and $\nu_{2}$.
Lastly, we make an assumption regarding the existence of a bounded conditional mean operator for the model $(\operatorname{\mathbb{P}}_{g(\theta,x)})_{x\in\mathcal{X}}$, which was introduced in Alquier and Gerber, (2024). Define $\operatorname{\mathbb{P}}_{X}:=\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{P}}^{i}$ for $\operatorname{\mathbb{P}}^{i}$ as in (23). The conditional mean embedding operator is used to bound the MMD between $\operatorname{\mathbb{P}}_{\theta}:=\operatorname{\mathbb{P}}_{X}\operatorname{\mathbb{P}}_{g(\theta,\cdot)}$ as defined in (24) and ${\operatorname{\mathbb{P}}^{0}_{\theta}}^{\texttt{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Classical}}}=\operatorname{\mathbb{P}}^{0}_{X}\operatorname{\mathbb{P}}_{g(\theta,\cdot)}$ or ${\operatorname{\mathbb{P}}^{0}_{\theta}}^{\texttt{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}Berkson}}}=\operatorname{\mathbb{P}}^{0,n}_{X}\operatorname{\mathbb{P}}_{g(\theta,\cdot)}$, in terms of the MMD between $\operatorname{\mathbb{P}}_{X}$ and $\operatorname{\mathbb{P}}_{X}^{0}$ or $\operatorname{\mathbb{P}}_{X}$ and $\operatorname{\mathbb{P}}_{X}^{0,n}$ respectively. Under Assumptions 4.0.1 and 4.0.2, the conditional mean embedding operator of a probability measure $\operatorname{\mathbb{P}}=\operatorname{\mathbb{P}}_{X}\operatorname{\mathbb{P}}_{Y\>|\>\cdot}\in\mathcal{P}(\mathcal{Z})$ is a mapping $\mathcal{C}_{Y\>|\>X}:\mathcal{H}_{\mathcal{X}}\rightarrow\mathcal{H}_{\mathcal{Y}}$ such that $\mathcal{C}_{Y\>|\>X}k_{X}(x,\cdot)=\mu_{\operatorname{\mathbb{P}}_{Y\>|\>X=x}}$ where $\mu_{\operatorname{\mathbb{P}}_{Y\>|\>X=x}}$ is the kernel mean embedding of $\operatorname{\mathbb{P}}_{Y\>|\>X=x}$. Using this definition we make the following assumption:
@@
A more detailed analysis of kernel mean embeddings can be found in Klebanov et al., (2020). We refer the reader to Alquier and Gerber, (2024) for a thorough discussion on sufficient conditions for the existence of a bounded linear conditional mean operator. Table 1 summarises the target joint and marginal distributions alongside the equivalent joint model probability measures for each of the two types of ME.
Berkson Classical
Target distribution of $(X,Y)$ $\operatorname{\mathbb{P}}^{0}:=\operatorname{\mathbb{P}}^{0,n}_{X}\operatorname{\mathbb{P}}^{0}_{Y\>|\>x}$ $\operatorname{\mathbb{P}}^{0}:=\operatorname{\mathbb{P}}^{0}_{X}\operatorname{\mathbb{P}}^{0}_{Y\>|\>x}$
Joint model of $(X,Y)$ $\operatorname{\mathbb{P}}^{0}_{\theta}:=\operatorname{\mathbb{P}}^{0,n}_{X}\operatorname{\mathbb{P}}_{g(\theta,\cdot)}$ $\operatorname{\mathbb{P}}^{0}_{\theta}:=\operatorname{\mathbb{P}}^{0}_{X}\operatorname{\mathbb{P}}_{g(\theta,\cdot)}$
Joint model of $(X,Y)$ from $\mu$ $\operatorname{\mathbb{P}}_{\theta}:=\operatorname{\mathbb{P}}_{X}\operatorname{\mathbb{P}}_{g(\theta,\cdot)}$ $\operatorname{\mathbb{P}}_{\theta}:=\operatorname{\mathbb{P}}_{X}\operatorname{\mathbb{P}}_{g(\theta,\cdot)}$
Table 1: Summary of notation for for Berkson and Classical error models.
@@ 4.2 Main results
We now bound the generalization error of our method. Proofs for all theoretical results of this section are provided in Appendix B. Note that even when the functional relationship $g$ is known, $\operatorname{\mathbb{P}}_{g(\theta,\cdot)}$ might still be misspecified so the expected error is trivially lower bounded by $\inf_{\theta\in\Theta}\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}_{\theta}^{0})$. Although this type of model misspecification is not directly targeted in this work, we expect to naturally remain robust to such misspecifications due to the robust nature of minimum MMD estimators as studied in Briol et al., (2019); Chérief-Abdellatif and Alquier, (2022); Dellaporta et al., (2022).
@@
Let $k^{2}_{X}=k_{X}\otimes k_{X}$ and $\operatorname{\mathbb{P}}^{0}=\operatorname{\mathbb{P}}_{X}^{0,n}\operatorname{\mathbb{P}}_{Y\>|\>x}^{0}$. Then under Assumptions 1-4 and $\Lambda$ as in Assumption 4.0.4 we have:
@@
$\overbrace{\mathbb{E}_{(x,y)_{1:n}\sim\operatorname{\mathbb{P}}_{X}^{0}}\left[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}\left[\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0})\right]\right]-\inf_{\theta\in\Theta}\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}^{0}_{\theta})}^{\text{Generalisation error}}$
@@
$\quad\quad\leq\frac{4}{\sqrt{n}}+2(1+\Lambda)\left(\sqrt{\frac{1}{c+2}}+\underbrace{\frac{c}{c+1}\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N},\operatorname{\mathbb{F}}_{N}^{0})}_{\text{Priorspecification}}+\underbrace{\frac{1}{c+1}\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})}_{\text{ME deviation from 0}}\right).$
@@
where expectation under $\operatorname{\mathbb{P}}:=(\operatorname{\mathbb{P}}^{1},\dots,\operatorname{\mathbb{P}}^{n})\sim\mu$ denotes the expectation under:
$\operatorname{\mathbb{P}}^{i}\sim\mu_{i},\quad\forall i=1,\dots,n.$
@@
Recall that for any set $A\in\sigma_{\mathcal{X}}$, $\operatorname{\mathbb{P}}^{0}_{X\>|\>w_{i}}(A)=\operatorname{\mathbb{F}}_{N}^{0}(A_{w_{i}})$. As expected, the upper bound of the generalization error depends on the DP prior both through the concentration parameter $c$ and the quality of the prior $\operatorname{\mathbb{F}}_{N}$ with respect to $\operatorname{\mathbb{F}}_{N}^{0}$ as well as the deviation of the true ME from zero. Theorem 4.1 highlights an error trade-off based on the selection of $c$ and the prior beliefs about $X\>|\>W=w_{i}$, providing a mathematical interpretation of the prior elicitation discussion in Section 3.3. The three discussed cases can be seen as follows:
  • $c\rightarrow\infty$: this case corresponds to putting all the posterior weight in (6) to the prior centering measure $\operatorname{\mathbb{F}}_{N}$. As a result, the generalization error is upper bounded by $\frac{4}{\sqrt{n}}+2(1+\Lambda)\operatorname{\operatorname{MMD}}_{k_{X}^{2}}(\operatorname{\mathbb{F}}_{N},\operatorname{\mathbb{F}}_{N}^{0})$, and it depends on the quality of $\operatorname{\mathbb{F}}_{N}$.
  • $c=0$: the prior belief is that there is no ME since the posterior centering measure is concentrated on the observed data. In this case, the generalization error in Theorem 4.1 is upper bounded by $\frac{4}{\sqrt{n}}+2(1+\Lambda)\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N},\delta_{0})+\frac{2(1+\Lambda)}{2^{1/2}}$ where the second term quantifies the distance of the true ME terms $\nu^{0}_{1:n}$ from zero.
  • $c=1$: The case of equally splitting the weight between the prior and observations in the nonparametric posterior centering measure. This can be seen in Theorem 4.1 since the two last terms are equally indexed by a constant.
@@
In the Classical error case, we get a similar bound which we present in Theorem 4.2. Unsurprisingly, the error bound now depends on the quality of the prior centering distribution with respect to $\operatorname{\mathbb{P}}_{X}^{0}$.
@@
Let $k^{2}_{X}=k_{X}\otimes k_{X}$ and $\operatorname{\mathbb{P}}^{0}=\operatorname{\mathbb{P}}_{X}^{0}\operatorname{\mathbb{P}}_{Y\>|\>X}^{0}$. Then under Assumptions 1-4 and $\Lambda$ as in Assumption 4.0.4 we have:
@@
$\quad\quad\leq\frac{4}{\sqrt{n}}+2(1+\Lambda)\left(\frac{1}{\sqrt{n}(c+1)}+\sqrt{\frac{1}{c+2}}+\underbrace{\frac{c}{c+1}\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{F}})}_{\text{Prior specification}}+\underbrace{\frac{1}{c+1}\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})}_{\text{ME deviation from 0}}\right).$
where $\operatorname{\mathbb{F}}:=\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{F}}_{X\>|\>w_{i}}$ and expectation under $\operatorname{\mathbb{P}}:=(\operatorname{\mathbb{P}}^{1},\dots,\operatorname{\mathbb{P}}^{n})\sim\mu$ denotes the expectation under:
$\operatorname{\mathbb{P}}^{i}\sim\mu_{i},\quad\forall i=1,\dots,n.$
Similar to the Berkson case, Theorem 4.2 provides clarity on the role of the prior specification. Here, $\operatorname{\mathbb{F}}$ incorporates prior beliefs about both the marginal distribution of $X$ through $\operatorname{\mathbb{F}}_{X}$ and about the ME distribution through $\operatorname{\mathbb{F}}_{W\>|\>X}$. Hence, the quality of the prior will be influenced by both these terms and will depend on how well the resulting measures $\operatorname{\mathbb{F}}_{X\>|\>w_{i}}$ describe the true distribution $\operatorname{\mathbb{P}}_{X}^{0}$.

### v3

@@ 4.1 Target Distribution & Assumptions
The kernel $k:\mathcal{Z}\times\mathcal{Z}\rightarrow\operatorname{\mathbb{R}}$ is $\mathfrak{S}_{\mathcal{Z}}$-measurable and satisfies $|k(z,z^{\prime})|\leq 1$ for any $z,z^{\prime}\in\mathcal{Z}$, where $|\cdot|$ denotes the euclidean norm . Moreover, $\phi:\mathcal{Z}\rightarrow\mathcal{H}_{k}$ is such that $k((x,y),(x^{\prime},y^{\prime}))=\left<\phi(x,y),\phi(x^{\prime},y^{\prime})\right>_{\mathcal{H}_{k}}$.
@@
There is a continuous kernel $k_{X}:\mathcal{X}\times\mathcal{X}\rightarrow\operatorname{\mathbb{R}}$ and a kernel $k_{Y}:\mathcal{Y}\times\mathcal{Y}\rightarrow\operatorname{\mathbb{R}}$ such that $|k_{X}(x,x^{\prime})|\leq 1$ for all $x,x^{\prime}\in\mathcal{X}$, $|k_{Y}(y,y^{\prime})|\leq 1$ for all $y,y^{\prime}\in\mathcal{Y}$ and $k=k_{X}\otimes k_{Y}$.
As noted in Alquier and Gerber, (2024), from Theorems 3-4 in Szabó and Sriperumbudur, (2017) it follows that under Assumption 4.0.2, $k$ is characteristic if $k_{X}$ and $k_{Y}$ are characteristic, continuous, translation invariant and bounded. When $\mathcal{X}=\operatorname{\mathbb{R}}^{d_{x}}$ and $\mathcal{Y}=\operatorname{\mathbb{R}}^{d_{y}}$ examples of such kernels $k_{X}$ and $k_{Y}$ include the widely-used RBF, exponential kernel and Matérn kernels. Here we restrict our results to the case where $k_{X}$ is a translation-invariant kernel:
@@
There exists a bounded, continuous, real-valued, positive-definite function $\psi$ on $\mathbb{R}^{d_{x}}$ such that $k_{X}(x,x^{\prime})=\psi(x-x^{\prime})$ for all $x,x^{\prime}\in\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}$. Here positive-definite means that for every points $x_{1},\dots,x_{n}\in\mathbb{R}^{d_{\mathcal{X}}}$, and coefficients $c_{1},\dots,c_{n}\in\mathbb{R}$,
$\sum_{i=1}^{n}\sum_{j=1}^{n}c_{i}c_{j}\,\psi(x_{i}-x_{j})\;\geq\;0.$
While this assumption is not strictly necessary for deriving generalisation error bounds, it demonstrates the error decomposition relative to the MMD between the true ME distribution and the prior centering measure. This choice also makes intuitive sense for the case of additive ME, since under a translation invariant kernel the distance between any two points $x+\nu_{1}$ and $x+\nu_{2}$ is characterized by the distance between $\nu_{1}$ and $\nu_{2}$.
Lastly, we make an assumption about the existence of a bounded conditional mean operator for the model $(\operatorname{\mathbb{P}}_{g(\theta,x)})_{x\in\mathcal{X}}$, which was introduced in Alquier and Gerber, (2024). Define $\operatorname{\mathbb{P}}_{X}:=\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{P}}^{i}$ for $\operatorname{\mathbb{P}}^{i}$ as in (17). The conditional mean embedding operator is used to bound the MMD between $\operatorname{\mathbb{P}}_{\theta}:=\operatorname{\mathbb{P}}_{X}\times\operatorname{\mathbb{P}}_{g(\theta,\cdot)}$ as defined in (18) and ${\operatorname{\mathbb{P}}^{0}_{\theta}}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\texttt{Classical}}}=\operatorname{\mathbb{P}}^{0}_{X}\times\operatorname{\mathbb{P}}_{g(\theta,\cdot)}$ or ${\operatorname{\mathbb{P}}^{0}_{\theta}}^{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\texttt{Berkson}}}=\operatorname{\mathbb{P}}^{0,n}_{X}\times\operatorname{\mathbb{P}}_{g(\theta,\cdot)}$, in terms of the MMD between $\operatorname{\mathbb{P}}_{X}$ and $\operatorname{\mathbb{P}}_{X}^{0}$ or $\operatorname{\mathbb{P}}_{X}$ and $\operatorname{\mathbb{P}}_{X}^{0,n}$ respectively. Under Assumptions 4.0.1 and 4.0.2, the conditional mean embedding operator of a probability measure $\operatorname{\mathbb{P}}=\operatorname{\mathbb{P}}_{X}\times\operatorname{\mathbb{P}}_{Y\>|\>X}\in\mathcal{P}_{\mathcal{Z}}$ is a mapping $\mathcal{C}_{Y\>|\>X}:\operatorname{\mathcal{H}_{k_{X}}}\rightarrow\operatorname{\mathcal{H}_{k_{Y}}}$ such that $\mathcal{C}_{Y\>|\>X}k_{X}(x,\cdot)=\mu_{\operatorname{\mathbb{P}}_{Y\>|\>X=x}}$ where $\mu_{\operatorname{\mathbb{P}}_{Y\>|\>X=x}}$ is the kernel mean embedding of $\operatorname{\mathbb{P}}_{Y\>|\>X=x}$. Using this definition we make the following assumption:
@@
A more detailed analysis of kernel mean embeddings can be found in Klebanov et al., (2020). We refer the reader to Alquier and Gerber, (2024) for a thorough discussion on sufficient conditions for the existence of a bounded linear conditional mean operator. Table 1 summarises the notation for the target joint and marginal distributions alongside the equivalent joint model probability measures for each of the two types of ME.
@@ 4.2 Main results
We now bound the generalisation error of our method. Proofs for all theoretical results of this section are provided in Appendix D. Note that even when the functional relationship $g$ is known, $\operatorname{\mathbb{P}}_{g(\theta,\cdot)}$ might still be misspecified so the expected error is trivially lower bounded by $\inf_{\theta\in\Theta}\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}_{\theta}^{0})$. Although this type of model misspecification is not directly targeted in this work, we expect to naturally remain robust to such misspecifications due to the robust nature of minimum MMD estimators as studied in Briol et al., (2019); Chérief-Abdellatif and Alquier, (2022); Dellaporta et al., (2022).
@@
Let $k^{2}_{X}=k_{X}\otimes k_{X}$ and $\operatorname{\mathbb{P}}^{0}=\operatorname{\mathbb{P}}_{X}^{0,n}\times\operatorname{\mathbb{P}}_{Y\>|\>X}^{0}$. Then under Assumptions 4.0.1-4.0.4 and $\Lambda$ as in Assumption 4.0.4 we have:
@@
$\overbrace{\mathbb{E}_{(x_{i},y_{i})_{1:n}\sim\operatorname{\mathbb{P}}^{0}}\left[\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}\left[\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0})\right]\right]-\inf_{\theta\in\Theta}\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}^{0}_{\theta})}^{\text{Generalisation error}}$
@@
$\quad\leq\frac{1}{\sqrt{n}}\left(2+\frac{2(1+\Lambda)}{\sqrt{c+2}}\right)+\underbrace{\frac{2c}{c+1}\left(\Lambda\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N},\operatorname{\mathbb{F}}_{N}^{0})+\sqrt{{\text{Var}}_{\mathcal{H}_{k_{X}}}[\operatorname{\mathbb{F}}_{N}^{0}]+\operatorname{\operatorname{MMD}}^{2}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\operatorname{\mathbb{F}}_{N})}\right)}_{\text{Prior specification \& ME RKHS variance}}$
$\quad\quad+\underbrace{\frac{2}{c+1}\left(\Lambda\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})+\sqrt{2}\sqrt{\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})}\right)}_{\text{ME deviation from 0}}.$
@@
where $\operatorname{\mathbb{P}}:=(\operatorname{\mathbb{P}}^{1},\dots,\operatorname{\mathbb{P}}^{n})\sim\mu$, $\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}$ denotes expectation under: $\{\operatorname{\mathbb{P}}^{i}\sim\mu_{i}\}_{1:n}$ and $\text{Var}_{\mathcal{H}_{k_{X}}}[\operatorname{\mathbb{F}}_{N}^{0}]$ denotes the RKHS variance of $\operatorname{\mathbb{F}}_{N}^{0}$, i.e. ${\text{Var}}_{\mathcal{H}_{k_{X}}}[\operatorname{\mathbb{F}}_{N}^{0}]:=\operatorname{\mathbb{E}}_{\nu\sim\operatorname{\mathbb{F}}_{N}^{0}}[k_{X}(\nu,\nu)]-\left<\mu_{\operatorname{\mathbb{F}}_{N}^{0}},\mu_{\operatorname{\mathbb{F}}_{N}^{0}}\right>_{\mathcal{H}_{k_{X}}}$.
@@
Recall that for any set $A\in\sigma_{\mathcal{X}}$, $\operatorname{\mathbb{P}}^{0}_{X\>|\>w_{i}}(A)=\operatorname{\mathbb{F}}_{N}^{0}(A_{w_{i}})$. As expected, the upper bound of the generalisation error depends on the DP prior both through the concentration parameter $c$ and the quality of the prior $\operatorname{\mathbb{F}}_{N}$ with respect to $\operatorname{\mathbb{F}}_{N}^{0}$ as well as the deviation of the true ME from zero and its RKHS variance. Theorem 4.1 highlights an error trade-off based on the selection of $c$ and the prior beliefs about $X\>|\>W=w_{i}$, providing a mathematical interpretation of the prior elicitation discussion in Section 3.3. The three discussed cases can be understood as follows: when $c\rightarrow\infty$, all of the posterior weight in (4) is placed on the prior centering measure $\operatorname{\mathbb{F}}_{N}$ and consequently the generalisation error upper bound depends on the quality of $\operatorname{\mathbb{F}}_{N}$ as well as a measure of dispersion of the true ME distribution. This is because if the ME is more dispersed, this leads to the distribution of $X\mid W$ being more dispersed, hence making the problem harder. When $c=0$ the posterior centring measure is concentrated on the observed data, hence the generalisation error upper bound depends on the $\sqrt{n}$ term and the last term, which quantifies the distance of the true ME terms $\nu^{0}_{1:n}$ from zero. Finally, $c=1$ corresponds to when the weight in (4) has been equally divided between $\operatorname{\mathbb{F}}_{N}$ and the observations. Similarly, in Theorem 4.1, the last two terms are equally weighted, and indexed by the same constant.
@@
In the Classical error case, we get a similar bound which we present in Theorem 4.2. Unsurprisingly, the error bound now depends on the quality of $\operatorname{\mathbb{F}}_{X\>|\>w_{i}}$ with respect to $\operatorname{\mathbb{P}}_{X}^{0}$.
@@
Let $k^{2}_{X}=k_{X}\otimes k_{X}$, $\operatorname{\mathbb{P}}^{0}=\operatorname{\mathbb{P}}_{X}^{0}\times\operatorname{\mathbb{P}}_{Y\>|\>X}^{0}$ and $\operatorname{\mathbb{F}}:=\frac{1}{n}\sum_{i=1}^{n}\operatorname{\mathbb{F}}_{X\>|\>w_{i}}$. Then under Assumptions 4.0.1-4.0.4 and $\Lambda$ as in Assumption 4.0.4 we have:
@@
$\leq\frac{1}{\sqrt{n}}\left(\frac{2\Lambda}{c+1}+\frac{2\Lambda+2}{\sqrt{c+2}}\right)$
$\quad+\underbrace{\frac{2c}{c+1}\left(\Lambda\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{F}})+\frac{1}{n}\sum_{i=1}^{n}\sqrt{{\text{Var}}_{\mathcal{H}_{k_{X}}}[\operatorname{\mathbb{P}}_{X}^{0}]+\operatorname{\operatorname{MMD}}^{2}_{k_{X}}(\operatorname{\mathbb{P}}_{X}^{0},\operatorname{\mathbb{F}}_{X\mid w_{i}})}\right)}_{\text{Prior specification \& RKHS variance of $X$}}$
$\quad+\underbrace{\frac{2}{c+1}\left(\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})+\sqrt{2}\sqrt{\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})}\right)}_{\text{ME deviation from 0}}.$
where $\operatorname{\mathbb{P}}:=(\operatorname{\mathbb{P}}^{1},\dots,\operatorname{\mathbb{P}}^{n})\sim\mu$ and $\operatorname{\mathbb{E}}_{\operatorname{\mathbb{P}}\sim\mu}$ denotes expectation under: $\{\operatorname{\mathbb{P}}^{i}\sim\mu_{i}\}_{1:n}$.
Similarly to the Berkson case, the bound on Theorem 4.2 depends on the prior specification. $\operatorname{\mathbb{F}}$ incorporates prior beliefs about both the marginal distribution of $X$ through $\operatorname{\mathbb{F}}_{X}$ and about the ME distribution through $\operatorname{\mathbb{F}}_{W\>|\>X}$. Hence, the quality of the prior depends on how well $\operatorname{\mathbb{F}}_{X\>|\>w_{i}}$ describes $\operatorname{\mathbb{P}}_{X}^{0}$. Moreover, similarly to the Berkson case, the prior specification part of the upper bound also depends on an RKHS variance term. In this case, the error depends on the variance of the marginal distribution of $X$. Intuitively, this is unsurprising since a larger dispersion of the distribution of $X$ leads to a larger variance of the distribution of $W\mid X$ hence making the estimation problem harder. This is more clearly illustrated in the special case of the Gaussian example presented in the next section.